# iIntel Labs, USA

### v1

@@ 6.4 Quantization Results for Object Detection
In Table 2 we present the quantization performance of OuroMamba and the baselines for object detection. Across different quantization settings and models, OuroMamba consistently outperforms the alternatives yielding up to $\mathbf{21.1}$ and $\mathbf{18.1}$ higher box AP and mask AP, respectively.
@@ 6.5 Quantization Results for Segmentation
6.6 Quantization Results for Diffusion Models
Quantitative Analysis. Following Hu et al. (2024), we assess the image generation fidelity of Zigma, a VMM-based diffusion model, under W4A4 quantization. Fidelity is measured using the Fréchet Inception Distance (FID) score Heusel et al. (2017) across 5k samples, with results summarized in Table 4. Notably, across both evaluated high-resolution datasets, the OuroMamba-quantized W4A4 Zigma model exhibits the smallest FID increase, maintaining high-quality image generation. In contrast, PTQ4VM, QMamba have significant degradation, highlighting their limitations in preserving fidelity under ultra low-precisions.
@@ 6.6 Quantization Results for Diffusion Models
  Faces HQ Landscape HQ
@@
Figure 6: Visualization of landscape Skorokhodov et al. (2021) image generation by Zigma under different W4A4 quantization techniques.
Qualitative Analysis. In Figure 6, we qualitatively compare the generated image quality of OuroMamba with the second best technique from Table 4, QMamba, for the landscape image generation quality Skorokhodov et al. (2021). As observed in Figure 6, OuroMamba generates images indistinguishable from the FP16 baseline, whereas QMamba produces significant noise artifacts and/or unreadable content.
@@ 6.7 GEMM Implementation and Evaluation
Figure 7: Overview of W4A4 hybrid GEMM kernel implementation.
@@
Implementation. Efficient inference of OuroMamba in VMMs requires addressing dynamic outlier extraction for real-time identification and mixed-precision GEMM operation for optimized computation. To this end, we implement the OuroMamba-Quant kernel using CUTLASS Thakkar et al. (2023). For efficient outlier channel extraction, activations are partitioned across channels and mapped to the thread blocks, where each block independently compares its assigned channels against the threshold $\theta$ to identify outliers and updates the $O_{\texttt{list}}$. To efficiently map the mixed-precision computation, we introduce a hybrid GEMM kernel. For a specific scenario of 4-bit weights with inlier and outlier activations of 4-bit and 8-bit, respectively, the hybrid GEMM executes one $4b\times 4b$ GEMM for inliers and one mixed-precision $4b\times 8b$ GEMM for outliers, as shown in Figure 7. We pack two consecutive 4-bit inlier activations into one byte, with outlier positions set to zero in the inlier buffer, and leverage the INT4 tensor cores for inlier GEMM. The 8-bit outlier channels are extracted and stored compactly in a separate contiguous outlier buffer. The outliers execute GEMM on INT8 tensor cores performing computations only on the combined tensor of extracted outlier columns. Finally, the outputs from the inlier and outlier GEMMs are dequantized to FP16 and summed together, with all steps fused into a single pipeline for efficiency.
@@
Evaluation. As shown in Figure 8, OuroMamba-Quant achieves speedup of $\mathbf{2.36}\times$ on Vim-B, $\mathbf{1.61}\times$ on VMamba-S and $\mathbf{1.63}\times$ on Zigma over FP16 baselines with a $\mathcal{B}=128$. The speedup primarily stems from optimizations performed for efficient outlier extraction, mixed-precision GEMM through the INT4 and INT8 tensor cores. Unlike PTQ4VM, that uses a standalone kernel for dequantization, we fuse dequantization directly into the GEMM pipeline, eliminating associated overhead and improving latency. Please refer appendix for additional details.
6.8 Discussions and Ablations
Effect of Batch Size $\mathcal{B}$. We present the accuracy comparison across batch sizes from 16 to 1024 in Figure 9(a) for W4A8 Vim-S on image classification. OuroMamba exhibits
Figure 8: End-to-end speedup over FP16 baseline.
minimal accuracy gains beyond 128, validating this as the optimal batch size. In contrast, VMM PTQ techniques PTQ4VM and QMamba rely on significantly larger batch sizes of 256 and 1024 real-data samples, respectively. Notably, even at batch sizes below 128, OuroMamba consistently outperforms PTQ4VM and QMamba, highlighting the effectiveness of its synthetic data generation which is tailored for quantization of the respective model.
Effect of Neighborhood Size $\mathcal{N}$ and Top-$n$ Patches. Figure 9(b) demonstrates the effect of different neighborhood sizes $\mathcal{N}$ and top-$n$ positive patch selection on the quality of synthetic data generated by OuroMamba-Gen. To assess this, we measure the top-1 accuracy of W4A8 Vim-S using the generated samples are as for calibration. Notably, a $5\times 5$ neighborhood size with the top-12 positive patches yields the highest accuracy. This choice of $\mathcal{N}$ effectively captures complex spatial dependencies while remaining computationally efficient compared to using all patches in $\alpha_{p}$.
Impact of Scanning Direction. To demonstrate that OuroMamba-Gen is scanning direction agnostic, we generate synthetic data by applying it independently to each scanning direction of VMamba-B. Using the same evaluation setup, we measure the top-1 accuracy for W4A4 quantization. As shown in Fig. fig. 9(c), under delta weighting, we observe that regardless of the scanning direction used for synthetic data generation, the resulting quantized model achieves similar accuracy.
Impact of Weighting Factor. Figure 9(c) shows the effect of uniform weighting compared to the $\Delta(t)$ weighting on the synthetic data generation by OuroMamba-Gen. For W4A4 VMamba-B, following the same evaluation setup as above, we see independent of different scanning directions delta weighting consistently yields accuracy improvements.
Choice of Objective Function. The study in Table 6 evaluates the impact of different loss function components in $\mathcal{L}^{gen}$ on synthetic data generation effectiveness and their influence on the top-1 accuracy of W4A8 quantization for Vim-S. We empirically validate our choice of patch-level contrastive learning ($\mathcal{L}^{C}$) over the patch-similarity metric ($\mathcal{L}^{PSE}$) used in prior ViT DFQ techniques Li et al. (2022; 2023b), demonstrating its superior effectiveness in generating high-quality calibration data. As shown in Table 6, a linear combination of $\mathcal{L}^{C}$ and $\mathcal{L}^{O}$ yields the highest accuracy, while using only $\mathcal{L}^{O}$ results in the lowest, highlighting its limited role in capturing semantic relationships for synthetic data generation. Although $\mathcal{L}^{PSE}$ combined with $\mathcal{L}^{O}$ achieves acceptable accuracy, it still underperforms compared to $\mathcal{L}^{gen}$.
Effect of $n_{\texttt{refresh}}$. In Figure 9(d), we present the normalized speedup for W4A4 Vim-B and VMamba-B across different values of the periodic refresh parameter $n_{\texttt{refresh}}$. We observe that $n_{\texttt{refresh}}=10$ time-steps yields the highest speedup. Reducing $n_{\texttt{refresh}}$ leads to a sharp decline due to frequent reinitialization of $O_{\texttt{list}}$ and repeated outlier extraction over the entire tensor. Conversely, increasing $n_{\texttt{refresh}}$ beyond 10 results in excessive outlier accumulation, increasing computational complexity and reducing speedup. Notably, when $n_{\texttt{refresh}}={Full}$ (i.e., no refresh), OuroMamba-Quant becomes slower than the FP16 baseline due to the overhead of processing accumulated outliers.
@@ 6.8 Discussions and Ablations
Table 6: Impact of different loss components.
@@
Limitation Discussion. The dynamic outlier selection assumes that inlier distributions remain stable relative to their values determined during calibration. In our experiments across all evaluated models, we observed no significant variations in the dynamic range of inliers. However, if future model architectures exhibit substantial fluctuations in inlier values, additional investigation will be necessary to adapt the selection process accordingly.
Appendix A Appendix
Appendix B Extended Related Works
Existing PTQ techniques for ViTs Lin et al. (2021); Jiang et al. (2025); Li et al. (2023c); Yang et al. (2024); Ma et al. (2024); Yuan et al. (2022) address long-tailed distributions and static activation outliers to enhance quantization accuracy. For instance, DopQ-ViT Yang et al. (2024) and ADFQ-ViT Jiang et al. (2025) mitigate outliers by optimizing per-channel and per-patch scale factors, respectively. FQ-ViT Lin et al. (2021) introduces Power-of-Two Factor (PTF) for inter-channel LayerNorm variation and Log-Int-Softmax (LIS) for 4-bit attention map quantization. Among the early PTQ methods for Mamba models, Mamba-PTQ Pierro & Abreu (2024) and Quamba Chiang et al. (2024) identified activation outliers as a key challenge but were tailored for language tasks. More recently, VMM PTQ techniques Cho et al. (2025); Li et al. (2025b) highlighted the highly dynamic activation distributions and inter-channel variations across time-steps. PTQ4VM Cho et al. (2025) adapts SmoothQuant Xiao et al. (2023) to migrate activation outliers into weights using a migration factor. However, as noted in Li et al. (2025a), this increases weight complexity, making both weights and activations more sensitive to dynamic variations, rendering it ineffective for ultra-low precision ($<4$ bits) quantization. Additionally, PTQ4VM does not quantize SSM operators, limiting its scope to linear layer weights and output activations. QMamba Li et al. (2025b) addresses the dynamic inter-time-step variations in VMMs’ hidden states by introducing fine-grained temporal grouped quantization, quantizing both weights and activations. Similarly, kSQ-VMM Shi et al. (2025) applies similarity-based k-scaled channel-wise and token-wise quantization to handle dynamic activation distributions. However, existing VMM PTQ methods rely on static scale factors Cho et al. (2025); Li et al. (2025b) or fixed temporal groupings Li et al. (2025b), leading to accuracy degradation at ultra-low bit precisions due to their inability to dynamically manage outlier channels at runtime.
Table 7: Quantization accuracy comparison of SoTA techniques on ImageNet classification. ‘R’, ‘S’ signifies real and synthetic calibration data.
  Method Data #Images W/A Top-1 W/A Top-1 W/A Top-1
  Vim-T Baseline - - 32/32 76.10 32/32 76.10 32/32 76.10
PTQ4VM Cho et al. (2025) R 256 4/8 74.15 6/6 73.94 4/4 56.29
QMamba Li et al. (2025b) R 1024 4/8 70.13 6/6 57.95 4/4 53.41
OuroMamba (Ours) S 128 4/8 74.98 6/6 74.84 4/4 63.49
       VMamba-T Baseline - - 32/32 82.60 32/32 82.60 32/32 82.60
PTQ4VM Cho et al. (2025) R 256 4/8 77.02 6/6 75.67 4/4 72.67
QMamba Li et al. (2025b) R 1024 4/8 76.51 6/6 80.49 4/4 51.48
OuroMamba (Ours) S 128 4/8 81.73 6/6 80.15 4/4 77.56
       Hybrid Model MambaVision-T Baseline - - 32/32 82.30 32/32 82.30 32/32 82.30
PTQ4VM Cho et al. (2025) R 256 4/8 72.13 6/6 69.39 4/4 67.67
QMamba Li et al. (2025b) R 1024 4/8 71.93 6/6 68.17 4/4 65.33
OuroMamba (Ours) S 128 4/8 80.57 6/6 79.05 4/4 74.92
Appendix C Additional Quantization Results
In Table 1 we provide additional quantization results of Vim-T Zhu et al. (2024), VMamba-T Liu et al. (2025) and the hybrid model MambaVision-T Hatamizadeh & Kautz (2025) for image classification.
Appendix D GEMM Implementation Details
We first describe how the GEMM operation can be decomposed into separate computations for outliers and inliers. Consider an output element $Y[i,j]$ computed as
$Y[i,j]=\sum_{k\in\mathcal{I}}A^{I}[i,k]\,W[k,j]+\sum_{k\in\mathcal{O}}A^{O}[i,k]\,W[k,j]$ (8)
where the inlier activations $A^{I}$ have the outlier positions zeroed out, and the outlier activations $A^{O}$ have the inlier positions zeroed. This decomposition guarantees that the sum of the two partial GEMM results yields the same $Y[i,j]$ as the original full GEMM.
We now introduce our GEMM pipeline, which consists of the following five steps:
Outlier Extraction. Outlier values in the input activation are identified, and their corresponding positions are zeroed out. The outlier columns are then compacted into a small INT8 outlier buffer.
Inlier Extraction. With the outlier positions already zeroed out, the inlier values are extracted and packed into INT4 buffers, storing two values per byte.
INT4 GEMM. An INT4 GEMM is performed on the inlier data. During the CUTLASS epilogue, the results are immediately dequantized by multiplying with the activation and weight scales. This fusion is enabled by the use of per-tensor quantization for inliers, which offers greater efficiency compared to the per-token inlier quantization employed by PTQ4VM.
INT8 GEMM. A mixed-input INT4-INT8 GEMM is executed between the inlier and the compacted outlier matrices, utilizing INT8 tensor cores.
Outlier Dequantization and Combination. Finally, the dequantization of outliers and the combination of the two GEMM results are fused into a single kernel. This kernel is memory-bound because it writes the final result matrix.
Figure 10: Quantization performance comparison for text-to-image generation task.
Appendix E Speed Breakdown Results
As shown in Figure 11, outlier extraction incurs minimal overhead. Specifically, we partition activation by channels, so outlier channel indices and scaling factors are calculated and recorded in parallel. Additionally, our compact extraction approach
Figure 11: Kernel breakdown of OuroMamba-Quant.
restricts the INT8 GEMM operation to a small subset of outlier channels, limiting its runtime contribution to less than 5%. As expected, the INT4 GEMM remains the dominant component. The primary performance bottleneck is the dequantization and combination step. This step writes to the entire output matrix, making it inherently memory-bound and therefore more expensive. Notably, the dequantization overhead is higher in Vmamba-B because it has a higher outlier rate (4.3%) compared to Vim-B (1.3%). Nonetheless, even in scenarios with higher outlier densities, our overall pipeline remains efficient due to the minimal costs associated with both outlier extraction and the outlier GEMM computations.
Appendix F Memory Compression Results
Table 8: Memory compression comparison over PTQ4VM Cho et al. (2025).
  Vim-S Vim-B
  Method W/A Mem. Comp. W/A Mem. Comp.
  Baseline 16/16 1.00 16/16 1.00
  PTQ4VM Cho et al. (2025) 4/4 1.81 4/4 2.02
OuroMamba (Ours) 4/4 3.63 4/4 3.80
As shown in Table 8, we compare the memory compression factor of OuroMamba with PTQ4VM at W4A4, using the FP16 model as the baseline, on the Vim-S and Vim-B models Zhu et al. (2024). The results show that OuroMamba consistently achieves a high memory compression factor of up to 3.80$\times$ compared to the baseline FP16 model, while PTQ4VM achieves a memory compression factor of only 2.02$\times$, as it quantizes only the Linear layers of VMMs.
Figure 12: Generated synthetic data samples.
Appendix G Extension of OuroMamba-Quant to Transformer based models
We extend OuroMamba-Gen to Transformer-based models and layers by mapping the time-step dimension to the token dimension. Outlier channels are identified per token, with $O_{\texttt{list}}$ propagated across tokens.
Appendix H Text-to-Image Generation Results
Implementation. We applied OuroMamba-Quant to PixArt-$\Sigma$ Chen et al. (2024a) with 20-iteration setting. Following ViDiT-Q Zhao et al. (2025), we quantize linear layers for query, key, and value projections and the second projection layer of feed-forward network to W4A4. Meanwhile, the first projection layer of feed-forward network and the output projection in self-attention are quantized in 8-bit for better numerical stability, while outliers bits are fixed at 8-bit and $n_{\texttt{refresh}}$ is set to 10. For calibration, we follow Q-Diffusion Li et al. (2023a) and randomly sample text prompts from MS-COCO dataset Lin et al. (2015) to obtain outlier threshold and inlier scale factors.
Results. In Figure 10, we visualize the generated images of W4A8, W4A4 Ouromamba-Quant quantized PixArt-$\Sigma$ compared to W4A8 Q-DiT Chen et al. (2024b) and W4A8 PTQ4DiT Wu et al. (2024).
Appendix I Additional Synthetic data samples
In Figure 12, we additionally visualize synthetic samples generated by OuroMamba-Gen for image classification, object detection and segmentation tasks for Vim-B model Zhu et al. (2024).

### v2

@@ 6.4 Quantization Results for Object Detection
Table 2 shows the quantization performance of OuroMamba and the baselines for object detection. Across different quantization settings, OuroMamba consistently yield superior performance yielding up to $\mathbf{21.1}$ and $\mathbf{18.1}$ higher box AP and mask AP, respectively.
@@ 6.5 Quantization Results for Segmentation
  Faces HQ [FFHQ_NVIDIA] Landscape HQ [skorokhodov2021aligning]
@@
6.6 Quantization Results for Diffusion Models
Quantitative Analysis. Following the experimental setup in [hu2024zigma], we assess the image generation fidelity of Zigma [hu2024zigma], a VMM-based diffusion model, under W4A4 quantization. Fidelity is measured using the Fréchet Inception Distance (FID) score [heusel2017gans] across 5k samples, with results summarized in Table 4. Notably, across both evaluated datasets, the OuroMamba-quantized W4A4 Zigma model exhibits the smallest FID increase, maintaining high image generation quality. PTQ4VM, QMamba suffer significant degradation, under ultra low-precisions.
Figure 5: Visualization of landscape [2021landscape] image generation by Zigma under different W4A4 quantization techniques.
Qualitative Analysis. In Fig. 5, we qualitatively compare the generated image quality of OuroMamba with the second best technique from Table 4, QMamba, for the landscape image generation quality [2021landscape]. As observed in Fig. 5, OuroMamba generates images indistinguishable from the FP16 baseline, whereas QMamba produces significant noise artifacts and/or unreadable content.
Figure 6: Overview of W4A4 hybrid GEMM kernel.
Figure 7: End-to-end latency speedup over FP16 baseline
@@ 6.7 GEMM Implementation and Evaluation
Implementation. Efficient inference of OuroMamba in VMMs requires addressing dynamic outlier extraction for real-time identification and mixed-precision GEMM operation for optimized computation. To this end, we implement the OuroMamba-Quant kernel using CUTLASS [cutlass]. For efficient outlier channel extraction, activations are partitioned across channels and mapped to the thread blocks, where each block independently compares its assigned channels against the threshold $\theta$ to identify outliers and updates the $O_{\texttt{list}}$. To efficiently map the mixed-precision computation, we introduce a hybrid GEMM kernel. For a specific scenario of 4-bit weights with inlier and outlier activations of 4-bit and 8-bit, respectively, the hybrid GEMM executes one $4b\times 4b$ GEMM for inliers and one mixed-precision $4b\times 8b$ GEMM for outliers, as shown in Fig. 6. We pack two consecutive 4-bit inlier activations into one byte, with outlier positions set to zero in the inlier buffer, and leverage the INT4 tensor cores for inlier GEMM. The 8-bit outlier channels are extracted and stored compactly in a separate contiguous outlier buffer. The outliers execute GEMM on INT8 tensor cores performing computations only on the combined tensor of extracted outlier columns. Finally, the outputs from the inlier and outlier GEMMs are dequantized to FP16 and summed together, with all steps fused into a single pipeline for efficiency.
@@
Evaluation. As shown in Fig. 7, OuroMamba-Quant achieves speedup of $\mathbf{2.36}\times$ on Vim-B, $\mathbf{1.61}\times$ on VMamba-S and $\mathbf{1.63}\times$ on Zigma over FP16 baselines with a $\mathcal{B}=128$. The speedup primarily stems from optimizations performed for efficient outlier extraction, mixed-precision GEMM through the INT4 and INT8 tensor cores. Unlike PTQ4VM, that uses a standalone kernel for dequantization, we fuse dequantization directly into the GEMM pipeline, eliminating associated overhead and improving latency. Please refer supplementary for additional details.
@@
Table 6: Impact of different loss components for data generation.
@@
6.8 Discussions and Ablations
Effect of Batch Size $\mathcal{B}$. We present the accuracy comparison across batch sizes from 16 to 1024 in Fig. 8(a) for W4A8 Vim-S on image classification. OuroMamba exhibits minimal accuracy gains beyond 128, validating this as the optimal batch size. In contrast, VMM PTQ techniques PTQ4VM and QMamba rely on significantly larger batch sizes of 256 and 1024 real-data samples, respectively. Notably, even at batch sizes below 128, OuroMamba consistently outperforms PTQ4VM and QMamba, highlighting the effectiveness of its synthetic data generation which is tailored for quantization of the respective model.
Effect of Neighborhood Size $\mathcal{N}$ and Top-$n$ Patches. Fig. 8(b) demonstrates the effect of different neighborhood sizes $\mathcal{N}$ and top-$n$ positive patch selection on the quality of synthetic data generated by OuroMamba-Gen. To assess this, we measure the top-1 accuracy of W4A8 Vim-S using the generated samples are as for calibration. Notably, a $5\times 5$ neighborhood size with the top-12 positive patches yields the highest accuracy. This choice of $\mathcal{N}$ effectively captures complex spatial dependencies while remaining computationally efficient compared to using all patches in $\alpha_{p}$.
Impact of Scanning Direction. We generate synthetic data by applying it independently to each scanning direction of VMamba-B. Using the same evaluation setup, we measure the top-1 accuracy for W4A4 quantization. As shown in Fig. Fig. 8(c), under delta weighting, we observe that regardless of the scanning direction, the resulting model achieves similar accuracy.
Impact of Weighting Factor. Fig. 8(c) shows the effect of uniform weighting compared to the $\Delta(t)$ weighting on the synthetic data generation by OuroMamba-Gen. For W4A4 VMamba-B, following the same evaluation setup as above, we see independent of different scanning directions delta weighting consistently yields accuracy improvements.
Choice of Objective Function. The study in Table 6 evaluates the impact of different loss function components in $\mathcal{L}^{gen}$ on synthetic data generation effectiveness and their influence on the top-1 accuracy of W4A8 quantization for Vim-S. We empirically validate our choice of patch-level contrastive learning ($\mathcal{L}^{C}$) over the patch-similarity metric ($\mathcal{L}^{PSE}$) used in prior ViT DFQ techniques [li2022patch, li2023psaq], demonstrating its superior effectiveness in generating high-quality calibration data. As shown in Table 6, a linear combination of $\mathcal{L}^{C}$ and $\mathcal{L}^{O}$ yields the highest accuracy, while using only $\mathcal{L}^{O}$ results in the lowest, highlighting its limited role in capturing semantic relationships for synthetic data generation. Although $\mathcal{L}^{PSE}$ combined with $\mathcal{L}^{O}$ achieves acceptable accuracy, it still underperforms compared to $\mathcal{L}^{gen}$.
Effect of $n_{\texttt{refresh}}$. In Fig. 8(d), we present the normalized speedup for W4A4 Vim-B and VMamba-B across different values of the periodic refresh parameter $n_{\texttt{refresh}}$. We observe that $n_{\texttt{refresh}}=10$ time-steps yields the highest speedup. Reducing $n_{\texttt{refresh}}$ leads to a sharp decline due to frequent reinitialization of $O_{\texttt{list}}$ and repeated outlier extraction over the entire tensor. Conversely, increasing $n_{\texttt{refresh}}$ beyond 10 results in excessive outlier accumulation, increasing computational complexity and reducing speedup. Notably, when $n_{\texttt{refresh}}={Full}$ (i.e., no refresh), OuroMamba-Quant becomes slower than the FP16 baseline due to the overhead of processing accumulated outliers.
@@ 6.8 Discussions and Ablations
Limitation Discussion. The dynamic outlier selection assumes that inlier distributions remain stable relative to their values determined during calibration. In our experiments across all models, we observed no significant variations in the dynamic range of inliers. However, if future model architectures exhibit substantial fluctuations in inlier values, additional investigation will be necessary.
7 Conclusions
This paper presents OuroMamba, a two-stage DFQ framework that addresses key challenges in enabling DFQ for VMMs. OuroMamba-Gen enhances implicit attention through patched neighborhood interactions and applies patch-level contrastive learning to generate semantically meaningful synthetic data. OuroMamba-Quant employs mixed-precision quantization with dynamic outlier detection, minimizing quantization error. Extensive evaluations demonstrate the SoTA quantization performance and speedup of our custom kernel across diverse vision tasks.
Acknowledgments
This work was supported in part by CoCoSys, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.